solve the ODE (3x2+4xy-6)dy+(6xy+2y2-5)=0


1
Expert's answer
2021-11-23T12:57:37-0500

 

"(6xy+2y^2-5)dx +(3x^2+4xy-6) dy=0"

 


"M(x,y)=6xy+2y^2-5"




"=> {\partial M(x,y)\over \partial y}=M_y=6x+4y"


"N(x,y)=3x^2+4xy-6"




"=> {\partial N(x,y)\over \partial x}=N_x=6x+4y"

Hence exact since "M_y=N_x"


There exist a function "u(x,y)" such that


"{\partial u(x,y) \over \partial x}=M(x,y)=6xy+2y^2-5"




"{\partial u(x,y) \over \partial y}=N(x,y)=3x^2+4xy-6"


Hence


"u(x, y)=\int(6xy+2y^2-5)dx+\varphi(y)"

"=3x^2y+2xy^2-5x+\varphi(y)"

Differentiating with respect to "y," we substitute the function into the second equation:


"{\partial u(x,y) \over \partial y}=3x^2+4xy+\varphi'(y)"

"=N(x,y)=3x^2+4xy-6"

Then


"\varphi'(y)=-6"

Integrate


"\varphi(y)=-\int 6dy=-6y-C"

"u(x, y)=3x^2y+2xy^2-5x-6y-C"

The general solution is

"3x^2y+2xy^2-5x-6y=C"